Paul invested $12,000 some at 14% annual interest and the rest at 10%. Last year he earned $1632 in interest. How much money was invested at each rate?
Thanks so much for any help.

Answer by jai_kos(139)

suppose $x is invested at 14% annual interest
interest after 1 year = $(14x)/100

rest $(12000-x)at 10% annual interest
interest after 1 year =$(12000-x)10/100

by the problem 14x/100 + (12000-x)10/100 =1632
or 14x/100 + 1200 -10x/100 =1632
or 4x/100=1632-1200=432
or x/25=432
or x=432*25= 10800
ans: $10800 is invested at 14%
$(12000-10800)=$1200 is invested at 10%